LIST SEPARATELY, if possible, two NEGATIVE INTEGERS such that one is double the other and their product is 98.

Answer by windsolace(8)

Let the smaller unknown number be x.
Therefore the larger number is 2x
x * 2x = 2(x^2)
2(x^2) = 98
x^2 = 98/2
x^2 = 49
x = 7
Since both integers are negative, therefore the 2 integers are -7 and -14
To test:
(-7) * (-14) = 98

I used * as multiply because x is already used as a constant in the solution. Don't want you people to get mixed up. =)